Two cars start 200 meters apart and drive twords each other car A has a constant speed of 5 m/s and car B has a constant speed of 10 m/s determine where and when they will meet


1
Expert's answer
2021-08-09T06:51:06-0400

The relative speed of the car A with respect to car B

"v=v_A+v_B=5\:\rm m/s+10\:\rm m/s=15\:\rm m/s"

Hence, the time of motion to the meet

"t=\frac{d}{v}=\frac{200\:\rm m}{15\:\rm m/s}=13.3\:\rm s"